# The Dirichlet heat trace for domains with curved corners

### v1

@@ 3.1. The conformal model domain
  • $\kappa_{1}$ (on the image of the $\theta=0$ ray): $2\Im(z_{0})$
  • $\kappa_{2}$ (on the image of the $\theta=\alpha$ ray): $2\Im(z_{0}e^{i\alpha})$
Here the orientation is chosen so that the signed curvatures measure curvature in the inward-pointing direction, so that a convex domain will have $\kappa_{1}\geq 0$ and $\kappa_{2}\geq 0$. A domain with symmetry about $\arg(z)=\alpha/2$ will have $\kappa_{1}=\kappa_{2}$.
@@
$\kappa_{-}$ $=-2\Im((u_{0}+iv_{0})(\cos\alpha_{0}+i\sin\alpha_{0}))=-2(u_{0}\sin\alpha_{0}+v_{0}\cos\alpha_{0})$
@@
(11) $\begin{cases}u_{0}=&-\frac{1}{2}\kappa_{+}\cot\alpha+\frac{1}{2}\kappa_{-}\csc\alpha;\\ v_{0}=&\frac{1}{2}\kappa_{+}.\end{cases}$
@@
(12) $r_{0}=\frac{1}{2}\csc\alpha\sqrt{\kappa_{+}^{2}-2\kappa_{+}\kappa_{-}\cos\alpha+\kappa_{-}^{2}};\quad\theta_{0}=\textrm{arg}(u_{0}+iv_{0}).$
@@
We now construct the heat kernel for $\mathcal{L}$. We use the techniques of [NRS], working at td, then at sf and ff. This corresponds to building a parametrix for the heat kernel, first at the interior diagonal, and then at the diagonal both near the boundary and near a corner. Once we have a sufficiently accurate parametrix, we iterate away the error via a Volterra series to produce the true heat kernel.
@@ 3.2. The heat kernel at td
It is proved in [NRS] that the heat kernel for $\mathcal{L}$, being the pullback of the heat kernel for the Laplacian on the curvilinear polygon $\Omega$, is polyhomogeneous conormal on the double space of [NRS]. We also know that its full expansion at td is the same as that for the heat kernel of the operator $\Phi^{*}(\Delta_{\mathbb{R}^{2}})=F^{-1}\Delta_{\mathbb{R}^{2}}$. This expansion may be written in terms of the (interior) boundary defining function $T=\sqrt{t}$ as
@@
We now use the work of Grieser [Grieser] to compute $a_{-2}$ and $a_{-1}$. We use his notation throughout this subsection: the coordinates are $(t,x,y)$, where $x\in\mathbb{R}^{2}$ is an interior coordinate on the first factor and $y\in\mathbb{R}^{2}$ on the second. The initial parametrix for the heat kernel is
(13) $K_{1}(t,x,y)=\frac{1}{4\pi t}\exp(-\frac{F(y)|x-y|^{2}}{4t}).$
@@
(14) $R_{1}(t,x,y):=(\partial_{t}-F^{-1}(x)\Delta_{\mathbb{R}^{2}})K_{1}(t,x,y)=(1-\frac{F(y)}{F(x)})\frac{\partial}{\partial t}K_{1}(t,x,y).$
The prefactor is $O(\sqrt{t})$ near the diagonal, so $R_{1}$ is $t^{1/2}$ less singular than $\partial_{t}K_{1}$.
@@
(15) $A(t,x,y)\sim\sum_{q\in d+\frac{1}{2}\mathbb{N}_{0}}t^{q}\,\Phi_{q}(A)(X,y)$
as $t\downarrow 0$ with $X$ fixed. The coefficient $\Phi_{q}(A)$ is the td-symbol of order $q$ of the operator $A$; if $A$ has order $d$, then $\Phi_{d}(A)$ is the principal td-symbol of $A$. Referring to Grieser [Grieser] for the definition of the various operator spaces $\Psi_{H}^{d}$, we have $K_{1}(t,x,y)\in\Psi_{H}^{-1}$. Roughly, this means that $K_{1}(t,x,y)$ has an expansion of the form (15) with d=1d=-1, and with coefficients decaying to infinite order at $|X|=\infty$.
To find the principal td-symbol of $K_{1}$, we compute, with $X=(x-y)/\sqrt{t}$,
$K_{1}(t,X\sqrt{t}+y,y)=\frac{1}{4\pi t}\exp(-\frac{F(y)}{4}X^{2}),$
so
(16) $\Phi_{-1}(K_{1})(X,y)=\frac{1}{4\pi}\exp(-\frac{F(y)}{4}X^{2}).$
We also analyze $R_{1}(t,X,y)$. Writing it out we obtain
(17) $R_{1}(t,X,y)=(\frac{F(X\sqrt{t}+y)-F(y)}{F(X\sqrt{t}+y)})(\frac{F(y)}{4}X^{2}-1)\frac{1}{t}K_{1}(t,X,y).$
The leading order term of this expression at $t=0$ is
(18) $(\frac{\sqrt{t}X\cdot\nabla(F)(y)}{F(y)})(\frac{F(y)}{4}X^{2}-1)\frac{1}{t}K_{1}(t,X,y).$
Therefore $R_{1}\in\Psi_{H}^{-1/2}$ with
(19) $\Phi_{-1/2}(R_{1})(X,y)=(\frac{X\cdot\nabla(F)(y)}{F(y)})(\frac{F(y)}{4}X^{2}-1)\frac{1}{4\pi}\exp(-\frac{F(y)}{4}X^{2}).$
By the heat kernel construction in Grieser [Grieser], the leading and subleading terms of the true heat kernel at td will be the same as those of
@@
$K_{1}$ has zero subleading term in the coordinates $(t,X,y)$, so for the subleading term we just need to understand $K_{1}*R_{1}$, in particular just the leading term of $K_{1}*R_{1}$, which is $\Phi_{-3/2}(K_{1}*R_{1})$. This can be computed using [Grieser, Proposition 2.6]. Observe that the integration in $Z$ is with respect to F(y)dZF(y)\,dZ, the metric at $y$. We obtain, including the minus sign,
@@ 3.3. A parametrix for the heat kernel
where $H_{cone}$ is the Dirichlet heat kernel for an exact sector of angle $\alpha$, and where $G$ is a scalar function chosen as a corrector to allow $H^{(1)}$ to satisfy the heat equation to two orders at td. Specifically,
$G:=(1-\frac{1}{8}TX^{2}X\cdot\nabla F(y))\exp[-\frac{F(y)-1}{4}X^{2}],$
which, when multiplied by the Euclidean heat kernel (the model for $H_{cone}$ at td), yields (25). We do not claim that $G$ is polyhomogeneous on the double space; it may blow up at td depending on the sign of F(y)1F(y)-1 (because $X=|z-z^{\prime}|/T$). However, $G$ is polyhomogeneous away from td and $G\cdot H_{cone}$ is polyhomogeneous everywhere.
Now examine $G$ at ff, the corner front face. This face is the most singular stratum of the heat space, where $t\to 0$ and both points $x,y=0$. The function $G$ is equal to 1 at ff because $T=0$ at ff, and $F=1$ at ff. As a consequence, $H^{(1)}$ is in the kernel of the heat operator $\mathcal{L}$ to leading order at ff; we say that it “solves the model problem” to one order at ff. It also solves the model problem to two orders at td. And since $T=0$ at sf and $\exp[-\frac{1}{4}(F(y)-1)X^{2}]H_{cone}$ agrees to first order at sf with the initial parametrix $\exp[-\frac{1}{4}(F(y)-1)X^{2}]H_{\mathbb{R}^{2}_{+}}$ used by Grieser at sf [Grieser], $H^{(1)}$ also solves the model problem to one order at sf.
@@
In order to use the composition theorem of [NRS, Theorem 3.16], which requires one of the two operators to vanish to infinite order at td, we work with a slightly improved parametrix $H^{(2)}$ given by solving away the full expansion at td. This equals $H^{(1)}$ plus a term which is order zero at td, ff, and sf. In particular, we note that if we have
@@
then $R^{(2)}-R^{(1)}$ has order 2-2 at ff. $H^{(2)}-H^{(1)}$ also has order 2-2 at ff. Moreover, $R^{(2)}$ vanishes to infinite order at td and it has leading order 3-3 at both ff and sf.
@@
We are interested in finding its term of order 1-1 at ff (the sub-leading term). By the composition formula in [NRS], using in particular that
@@
This in turn [^2] The justification for using the simpler parametrix $H^{(1)}$ to compute the subleading term at the ff face is done by an analysis of operator orders within the composition calculus of [NRS]. Let $D=H^{(2)}-H^{(1)}$ be the correction term that improves the parametrix at the td face. By construction, the leading order of $D$ at the ff face is at least $T^{0}$. The difference between the first two terms of the respective Neumann series is $\Delta H=D-H^{(1)}*\mathcal{L}(D)-D*R^{(1)}-D*\mathcal{L}(D)$. A direct application of the composition formula shows that each term in this expression has a leading order of $T^{0}$ or higher at the ff face. For instance, the most singular term, $H^{(1)}*\mathcal{L}(D)$, has an order of at least 2+(2)+4=0-2+(-2)+4=0. Since the entire difference $\Delta H$ is no more singular than $T^{0}$, the expressions $H^{(1)}-H^{(1)}*R^{(1)}$ and $H^{(2)}-H^{(2)}*R^{(2)}$ must agree for all terms of order $T^{-1}$ and below. We are therefore justified in using the simpler expression to compute the desired subleading coefficient. is the same as that of
@@
A critical fact here is that by the composition formula in [NRS], the leading term of $H^{(1)}*R^{(1)}$ at ff depends only on the leading terms of $H^{(1)}$ and $R^{(1)}$ at ff, not on their leading terms at other boundary faces (in particular at sf). This is helpful for the following reasons.
@@ 3.4. Computations with $G$
With this notation, natural local coordinates near the interior of the front face ff are $(T,R,\theta,R^{\prime},\theta^{\prime})$, where $T$ is a boundary defining function. In these coordinates, we compute the first two terms in the expansion (28), of orders $T^{-2}$ and $T^{-1}$. To do this we compute two terms in the expansion of $GH_{cone}$. We also compute the leading $T^{-2}$ term in the expansion of $H_{cone}$ and the leading $T^{-3}$ term in the expansion of the second convolution factor, as by the composition formula in [NRS] their convolution will give a term of order $T^{-2-3+4}=T^{-1}$ at ff. To this end, we denote
@@
First, we have a formula for $H_{cone}$ due to Cheeger [Cheeger1983]:
$H_{cone}=\frac{1}{2t}\exp[-\frac{1}{4t}(R^{2}+(R^{\prime})^{2})]\sum_{j=1}^{\infty}I_{\mu_{j}}(\frac{RR^{\prime}}{2t})\varphi_{j}(\theta)\varphi_{j}(\theta^{\prime}),$
where $\varphi_{j}(\theta)$ is the $L^{2}$-normalized eigenfunction
@@
Since by scaling we have
$H_{cone}=T^{-2}a(R,\theta,R^{\prime},\theta^{\prime}),$
we conclude that
@@
(35) $G(T,R,\theta,R^{\prime},\theta^{\prime})=\Bigg[1-\frac{1}{8}TD^{2}\langle R\cos\theta-R^{\prime}\cos\theta^{\prime},R\sin\theta-R^{\prime}\sin\theta^{\prime}\rangle\cdot\nabla F(R^{\prime}T,\theta^{\prime})\Bigg]\\ \exp[-\frac{1}{4}(F(R^{\prime}T,\theta^{\prime}))D^{2}].$
@@
$\nabla G=-\frac{a_{0}}{2}\nabla_{R,\theta}(D^{2}S)+O(T);\quad\Delta G=-\frac{a_{0}}{2}T^{-1}\Delta_{R,\theta}(D^{2}S)+O(1).$

### v2

@@ 3.1. The conformal model domain
  • $\kappa_{+}$ (on the image of the $\theta=0$ ray): $2\Im(z_{0})$
  • $\kappa_{-}$ (on the image of the $\theta=\alpha$ ray): $2\Im(z_{0}e^{i\alpha})$
Here the orientation is chosen so that the signed curvatures measure curvature in the inward-pointing direction, so that a convex domain will have $\kappa_{+}\geq 0$ and $\kappa_{-}\geq 0$. A domain with symmetry about $\arg(z)=\alpha/2$ will have $\kappa_{+}=\kappa_{-}$.
@@
$\kappa_{-}$ $=-2\Im((u_{0}+iv_{0})(\cos\alpha+i\sin\alpha))=-2(u_{0}\sin\alpha+v_{0}\cos\alpha)$
@@
(25) $\begin{cases}u_{0}=&-\frac{1}{2}\kappa_{+}\cot\alpha-\frac{1}{2}\kappa_{-}\csc\alpha;\\ v_{0}=&\frac{1}{2}\kappa_{+}.\end{cases}$
@@
(26) $r_{0}=\frac{1}{2}|\csc\alpha|\sqrt{\kappa_{+}^{2}+2\kappa_{+}\kappa_{-}\cos\alpha+\kappa_{-}^{2}};\quad\theta_{0}=\textrm{arg}(u_{0}+iv_{0}).$
@@
We now construct the heat kernel for $\mathcal{L}$. We use the techniques of [10], working at td, then at sf and ff. This corresponds to building a parametrix for the heat kernel, first at the interior diagonal, and then at the diagonal both near the boundary and near a corner. Once we have a sufficiently accurate parametrix, we iterate away the error via a Volterra series to produce the true heat kernel.
@@ 3.2. The heat kernel at td
It is proved in [10] that the heat kernel for $\mathcal{L}$, being the pullback of the heat kernel for the Laplacian on the curvilinear polygon $\Omega$, is polyhomogeneous conormal on the double space of [10]. We also know that its full expansion at td is the same as that for the heat kernel of the operator $\Phi^{*}(\Delta_{\mathbb{R}^{2}})=F^{-1}\Delta_{\mathbb{R}^{2}}$. This expansion may be written in terms of the (interior) boundary defining function $T=\sqrt{t}$ as
@@
We now use the work of Grieser [6] to compute $a_{-2}$ and $a_{-1}$. We use his notation throughout this subsection: the coordinates are $(t,x,y)$, where $x\in\mathbb{R}^{2}$ is an interior coordinate on the first factor and $y\in\mathbb{R}^{2}$ on the second. The initial parametrix for the heat kernel is
(27) $K_{1}(t,x,y)=\frac{1}{4\pi t}\exp\left(-\frac{F(y)|x-y|^{2}}{4t}\right).$
@@
(28) $R_{1}(t,x,y):=(\partial_{t}-F^{-1}(x)\Delta_{\mathbb{R}^{2}})K_{1}(t,x,y)=\left(1-\frac{F(y)}{F(x)}\right)\frac{\partial}{\partial t}K_{1}(t,x,y).$
@@
(29) $A(t,x,y)\sim\sum_{q\in d-\frac{1}{2}\mathbb{N}_{0}}t^{-2-q}\,\Phi_{q}(A)(X,y)$
as $t\downarrow 0$ with $X$ fixed. The coefficient $\Phi_{q}(A)$ is the td-symbol of order $q$ of the operator $A$; if $A$ has order $d$, then $\Phi_{d}(A)$ is the principal td-symbol of $A$. Referring to Grieser [6] for the definition of the various operator spaces $\Psi_{H}^{d}$, we have $K_{1}(t,x,y)\in\Psi_{H}^{-1}$. Roughly, this means that $K_{1}(t,x,y)$ has an expansion of the form (29) with d=1d=-1, hence leading term of order $t^{-1}$, and with coefficients decaying to infinite order at $|X|=\infty$.
The prefactor is $O(\sqrt{t})$ near the diagonal. Since $K_{1}\in\Psi_{H}^{-1}$, one has $\partial_{t}K_{1}\in\Psi_{H}^{0}$, which has size $t^{-2}$, so $R_{1}(t,y+\sqrt{t}\,X,y)$ has size $t^{-3/2}$ in td coordinates. Therefore $R_{1}\in\Psi_{H}^{-1/2}$ with principal td-symbol
(30) $\Phi_{-3/2}(R_{1})(X,y)=(\frac{X\cdot\nabla(F)(y)}{F(y)})(\frac{F(y)}{4}X^{2}-1)\frac{1}{4\pi}\exp(-\frac{F(y)}{4}X^{2}).$
By the heat kernel construction in Grieser [6], the leading and subleading terms of the true heat kernel at td will be the same as those of
@@
$K_{1}$ has zero subleading term in the coordinates $(t,X,y)$, so for the subleading term we just need to understand $K_{1}*R_{1}$, in particular just the leading term of $K_{1}*R_{1}$, which is $\Phi_{-3/2}(K_{1}*R_{1})$. This can be computed using [6, Proposition 2.6]. Observe that the integration in $Z$ is with respect to F(y)dZF(y)\,dZ, the metric at $y$. We obtain, including the minus sign,
@@ 3.3. A parametrix for the heat kernel
where $H_{cone}$ is the Dirichlet heat kernel for an exact sector of angle $\alpha$, and where $G$ is a scalar function chosen as a corrector to allow $H^{(1)}$ to satisfy the heat equation to two orders at td.
Notation. In this subsection we set $T=\sqrt{t}$ and use the rescaled difference vector
$X:=\frac{z-z^{\prime}}{T}\in\mathbb{R}^{2},\qquad|X|=\frac{|z-z^{\prime}|}{T}.$
We write $X\cdot\nabla F(z^{\prime})$ for the Euclidean dot product and $|X|^{2}=X\cdot X$.
Specifically, define
(37) $G:=\Bigl(1-\frac{1}{8}\,T\,|X|^{2}\,(X\cdot\nabla F(z^{\prime}))\Bigr)\exp\Bigl[-\frac{F(z^{\prime})-1}{4}\,|X|^{2}\Bigr].$
Then $G$ times the Euclidean heat kernel (the td model for $H_{\operatorname{cone}}$) yields (36). We do not claim that $G$ is polyhomogeneous on the double space; if $F(z^{\prime})<1$ then the factor $\exp\bigl(\frac{1-F(z^{\prime})}{4}|X|^{2}\bigr)$ grows as $|X|\to\infty$ (note that $|X|=|z-z^{\prime}|/T$). However, $G$ is polyhomogeneous away from $\mathrm{td}$, and $G\,H_{cone}$ is polyhomogeneous on the double space. Moreover, $G\,H_{cone}$ satisfies Dirichlet boundary conditions [^2] The Neumann problem is harder precisely because such a parametrix would not necessarily inherit Neumann boundary conditions. This can likely be worked around but we do not do so here., inheriting this from $H_{cone}$.
Now examine $G$ at ff, the corner front face. This face is the most singular stratum of the heat space, where $t\to 0$ and both points $x,y=0$. The function $G$ is equal to 1 at ff because $T=0$ at ff, and $F=1$ at ff. As a consequence, $H^{(1)}$ is in the kernel of the heat operator $\mathcal{L}$ to leading order at ff; we say that it “solves the model problem” to one order at ff. It also solves the model problem to two orders at td. And since $T=0$ at sf and $\exp[-\frac{1}{4}(F(y)-1)X^{2}]H_{cone}$ agrees to first order at sf with the initial parametrix $\exp[-\frac{1}{4}(F(y)-1)X^{2}]H_{\mathbb{R}^{2}_{+}}$ used by Grieser at sf [6], $H^{(1)}$ also solves the model problem to one order at sf.
@@
In order to use the composition theorem of [10, Theorem 3.16], which requires one of the two operators to vanish to infinite order at td, we work with a slightly improved parametrix $H^{(2)}$ given by solving away the full expansion at td. This equals $H^{(1)}$ plus a term which is order zero at td, ff, and sf. Let $D:=H^{(2)}-H^{(1)}$. Then $D$ has order $\geq 0$ at ff. Hence if we use the notation
@@
then $R^{(2)}-R^{(1)}=\mathcal{L}(D)$ has order $\geq-2$ at ff and sf. Moreover, $R^{(2)}$ vanishes to infinite order at td and it has leading order 3-3 at both ff and sf.
@@
We are interested in finding its term of order 1-1 at ff (the sub-leading term). By the composition formula in [10], using in particular that
@@
This in turn [^3] The justification for using the simpler parametrix $H^{(1)}$ to compute the subleading term at the ff face is done by an analysis of operator orders within the composition calculus of [10]. Let $D=H^{(2)}-H^{(1)}$ be the correction term that improves the parametrix at the td face. By construction, the leading order of $D$ at the ff face is at least $T^{0}$. The difference between the first two terms of the respective Neumann series is $\Delta H=D-H^{(1)}*\mathcal{L}(D)-D*R^{(1)}-D*\mathcal{L}(D)$. A direct application of the composition formula shows that each term in this expression has a leading order of $T^{0}$ or higher at the ff face. For instance, the most singular term, $H^{(1)}*\mathcal{L}(D)$, has an order of at least 2+(2)+4=0-2+(-2)+4=0. Since the entire difference $\Delta H$ is no more singular than $T^{0}$, the expressions $H^{(1)}-H^{(1)}*R^{(1)}$ and $H^{(2)}-H^{(2)}*R^{(2)}$ must agree for all terms of order $T^{-1}$ and below. We are therefore justified in using the simpler expression to compute the desired subleading coefficient. is the same as that of
@@
A critical fact here is that by the composition formula in [10], the leading term of $H^{(1)}*R^{(1)}$ at ff depends only on the leading terms of $H^{(1)}$ and $R^{(1)}$ at ff, not on their leading terms at other boundary faces (in particular at sf). This is helpful for the following reasons.
@@ 3.4. Computations with $G$
With this notation, natural local coordinates near the interior of the front face ff are $(T,R,\theta,R^{\prime},\theta^{\prime})$, where $T$ is a boundary defining function. In these coordinates, we compute the first two terms in the expansion (40), of orders $T^{-2}$ and $T^{-1}$. To do this we compute two terms in the expansion of $GH_{cone}$. We also compute the leading $T^{-2}$ term in the expansion of $H_{cone}$ and the leading $T^{-3}$ term in the expansion of the second convolution factor, as by the composition formula in [10] their convolution will give a term of order $T^{-2-3+4}=T^{-1}$ at ff. To this end, we denote
@@
Let $\varphi_{j}(\theta)$ be the $L^{2}$-normalized eigenfunction
@@
First, we use Cheeger’s formula for the Dirichlet heat kernel on the exact sector of angle $\alpha$ [2]. In polar coordinates $z=(r,\theta)$ and $z^{\prime}=(r^{\prime},\theta^{\prime})$, one has
$H_{\mathrm{cone}}(t;r,\theta,r^{\prime},\theta^{\prime})=\frac{1}{2t}\exp\Big[-\frac{1}{4t}\big(r^{2}+(r^{\prime})^{2}\big)\Big]\sum_{j=1}^{\infty}I_{\mu_{j}}\Big(\frac{rr^{\prime}}{2t}\Big)\,\varphi_{j}(\theta)\,\varphi_{j}(\theta^{\prime}).$
Now set $T=\sqrt{t}$, $R=r/T$, and $R^{\prime}=r^{\prime}/T$. Since $r=TR$, $r^{\prime}=TR^{\prime}$, and $t=T^{2}$, this becomes $H_{\mathrm{cone}}=T^{-2}a(R,\theta,R^{\prime},\theta^{\prime})$ by a simple scaling argument, with $a$ given by
@@
To be absolutely clear, we write
$H_{\mathrm{cone}}(t;r,\theta,r^{\prime},\theta^{\prime})=T^{-2}\,a\left(\frac{r}{\sqrt{t}},\theta,\frac{r^{\prime}}{\sqrt{t}},\theta^{\prime}\right),\qquad T=\sqrt{t},$
@@
(47) $G(T,R,\theta,R^{\prime},\theta^{\prime})=\Bigg[1-\frac{1}{8}TD^{2}\langle R\cos\theta-R^{\prime}\cos\theta^{\prime},R\sin\theta-R^{\prime}\sin\theta^{\prime}\rangle\cdot\nabla F(R^{\prime}T,\theta^{\prime})\Bigg]\\ \exp[-\frac{1}{4}(F(R^{\prime}T,\theta^{\prime})-1)D^{2}].$
@@
$\nabla G=-\frac{r_{0}}{2}\nabla_{R,\theta}(D^{2}S)+O(T);\quad\Delta G=-\frac{r_{0}}{2}T^{-1}\Delta_{R,\theta}(D^{2}S)+O(1).$